# Mexa: Multilingual Evaluation of English-Centric LLMs via Cross-Lingual Alignment

### v1

@@ Abstract
English-centric large language models (LLMs) often show strong multilingual capabilities. However, the multilingual performance of these models remains unclear and is not thoroughly evaluated for many languages. Most benchmarks for multilinguality focus on classic NLP tasks, or cover a minimal number of languages. We introduce Mexa, a method for assessing the multilingual capabilities of pre-trained English-centric LLMs using parallel sentences, which are available for more languages than existing downstream tasks. Mexa leverages the fact that English-centric LLMs use English as a kind of pivot language in their intermediate layers. It computes the alignment between English and non-English languages using parallel sentences to evaluate the transfer of language understanding from English to other languages. This alignment can be used to estimate model performance in other languages. We conduct studies using various parallel datasets (FLORES-200 and Bible), models (Llama family, Gemma family, Mistral, and OLMo), and established downstream tasks (Belebele, m-MMLU, and m-ARC). We explore different methods to compute embeddings in decoder-only models. Our results show that Mexa, in its default settings, achieves a statistically significant average Pearson correlation of 0.90 with three established downstream tasks across nine models and two parallel datasets. This suggests that Mexa is a reliable method for estimating the multilingual capabilities of English-centric LLMs, providing a clearer understanding of their multilingual potential and the inner workings of LLMs.
hf.co/spaces/cis-lmu/Mexa |
@@ 1 Introduction
Most state-of-the-art autoregressive large language models (LLMs) are English-centric, closed-source models such as Claude 3 Opus, GPT-4, and Gemini Pro (Anthropic, 2023; OpenAI et al., 2023; Gemini Team et al., 2023); open-weight models such as Llama 3.1, Gemma 2, and Mixtral (Dubey et al., 2024; Gemma Team et al., 2024b; Jiang et al., 2024); and open-source models such as OLMo (Groeneveld et al., 2024). Except for open-source models, where the data is available and thus the language distribution is transparent, there is confusion regarding the capabilities/language distribution of these LLMs in other languages.
@@
Primarily, the focus in evaluating LLMs has been on developing benchmarks to assess their performance in English. Most benchmarks in multilingual setups consist of classical monolingual NLP tasks such as sequence labeling (Ahuja et al., 2023; Lai et al., 2023a), the automatic translation of popular English benchmarks such as MMLU (Hendrycks et al., 2021) into a limited number of languages (Lai et al., 2023b; OpenAI, 2024), or language-specific benchmarks for languages such as Persian (Ghahroodi et al., 2024), Arabic (Koto et al., 2024), Korean (Son et al., 2024), Turkish (Yüksel et al., 2024), and Chinese (Li et al., 2024c).
@@
Most LLMs are English-centric, either by choice or due to the availability of abundant data sources in English. Either way, for these models to be effective in other languages, it is important that the other languages align with the main language, i.e., English. Given such alignment, English could act as a “rising tide that raises all ships,” meaning that improvements in English performance could benefit other languages, especially in tasks such as reasoning (Zhu et al., 2024). Contrarily, if a language does not align well with English, an English-centric LLM may not provide meaningful coverage for that language. Indeed, Wendler et al. (2024) have found that for Llama 2 (Touvron et al., 2023b), an English-centric LLM, English could be seen as a kind of “pivot” language, enabling to solve complex semantic tasks in a foreign language through a detour into English. More precisely, they show that Llama 2 was able to decode semantically correct next tokens in the middle layers, assigning higher probabilities to the English tokens than to the foreign version, which is only selected in the upper layers. Zhao et al. (2024) present a hypothesis regarding the middle layers of English-centric LLMs, suggesting that these models use English as a means of reasoning while incorporating multilingual knowledge. Based on their analysis, the number of language-specific neurons in the middle layers decreases within the self-attention mechanism but remains consistent across the layers of the feed-forward structure when processing multilingual queries.
@@
In this paper, we introduce Mexa, a method that uses the observation that English-centric LLMs semantically use English as a kind of pivot language in their middle layers to evaluate the actual multilingual coverage of LLMs. This is done by measuring how well the embeddings of parallel sentences in the middle layers of LLMs for non-English languages are aligned with the embedding of their corresponding English. We extensively verify the Mexa estimation of language coverage for each LLM, using Pearson correlation between estimated and actual scores for various tasks. We use two parallel datasets: FLORES-200 (NLLB Team et al., 2022) and Bible (Mayer & Cysouw, 2014); nine LLMs: Llama family, Gemma family, Mistral, and OLMo; and three tasks: Belebele (Bandarkar et al., 2024), m-MMLU, and m-ARC (Lai et al., 2023b). Our results show that Mexa achieves a promising average Pearson correlation of 0.90 with established downstream tasks across nine models and two parallel datasets. In our study on the calculation of Mexa scores, we conduct multiple design analyses to examine the impact of token-level pooling for the embeddings (i.e., using the last token versus a weighted average) and layer-level pooling in computing alignment scores. While Mexa demonstrates a high correlation across most setups, we find that a weighted average based on tokens, combined with mean pooling, yields the best results.
@@ 2 Background and Related Work
We discuss the distribution of pre-training data in LLMs and multilingual evaluation benchmarks in Appendices A.1 and A.2 while focusing on cross-lingual alignment here. Research in the cross-lingual alignment field either aims to uncover the underlying mechanisms of alignment and assess its impact on models and downstream tasks, or attempts to enhance model performance by enforcing alignment before, during, or after the pre-training phase. Most of these papers have focused on encoder-only models, such as XLM-R (Conneau et al., 2020a) and mBERT (Devlin et al., 2019), among others (Hämmerl et al., 2024). In this work, we focus primarily on decoder-only models.
@@
Understanding Alignment. Ye et al. (2023) show that English-centric models such as Llama 1 (Touvron et al., 2023a) not only possess multilingual transfer abilities (after fine-tuning on one source language, they can be applied to other languages) but may even surpass the multilingual transfer abilities of multilingual pre-trained models such as BLOOM (BigScience Workshop et al., 2023). Schäfer et al. (2024) find that GPT-2-style decoder-only models show strong cross-lingual generalization when trained on an imbalanced mix of languages. However, when trained on a balanced language set, they do not observe increased performance compared to monolingual settings. Wendler et al. (2024) perform single-token analysis to demonstrate that English-centered LLMs, such as Llama 2, use English semantically as an internal latent language in the middle layers when handling multilingual queries. Zhong et al. (2024) extend this analysis to multiple tokens, also showing that an LLM dominated by both English and Japanese uses both languages as internal latent languages. Zhao et al. (2024) explore how LLMs handle multilingualism. They hypothesize that LLMs initially interpret the query, converting multilingual inputs into English for task-solving. In the middle layers, the models rely on English with self-attention mechanisms for reasoning, while employing multilingual knowledge through feed-forward structures. In the final layers, LLMs generate responses consistent with the original query language. Li et al. (2024f) and Li et al. (2024b) are even more closely related to ours. Li et al. (2024f) employs absolute cosine similarity values between last token embeddings derived from parallel sentences with English to predict the ranking of language performance across various models. However, as we discuss in Section 3, relying solely on absolute cosine values can be misleading, and as shown in Section 5.3, absolute cosine values are less correlated with downstream tasks than Mexa score. Li et al. (2024b) uses English probing tasks and their automatic translations to construct a multilingual evaluation. While they compare embedding similarity scores between high- and low-resource languages with corresponding evaluation results, similar to Li et al. (2024f), they do not assess whether these correlations hold across other downstream tasks. In Section 5, we demonstrate that Mexa scores align closely with a broad range of downstream tasks.
@@
Boosting Alignment. The idea to enforce alignment in encoder-only models using parallel sentences dates back to Conneau & Lample (2019), and has been explored under various guises e.g., using mixed-language sentences and/or bilingual dictionaries (Huang et al., 2019; Conneau et al., 2020b; Cao et al., 2020; Kulshreshtha et al., 2020; Efimov et al., 2023; Zhang et al., 2023b). Recently, Li et al. (2024d) improve multilingual alignment by initializing the decoder-only models to generate similar representations of aligned words using contrastive learning and preserves this alignment using a code-switching strategy during pretraining. Liu et al. (2024a) propose a data allocation technique to select a core subset of languages for fine-tuning, better aligning the multilingual capabilities of decoder-only LLMs and making them more truthful in their responses. Li et al. (2024a) propose aligning internal sentence representations across different languages using multilingual contrastive learning and aligning outputs by following cross-lingual instructions in the target language for decoder-only models.
@@ 3 Mexa
We now describe the Mexa method for computing the alignment score of language with respect to the pivot language , given the language model . In this paper, we use the term cross-lingual alignment, geometric alignment, or simply alignment to refer to the semantic similarity of multilingual embeddings across languages. , for English-centric LLMs and in this paper, is English. To assess alignment, we use parallel sentences in two languages, and .
What defines semantic similarity in multilingual embeddings across languages? The goal of semantic similarity is to ensure that parallel sentences have sufficiently high similarity, reflecting alignment between the two languages. However, considering only the absolute cosine similarity value as the alignment score does not guarantee proper alignment. For some languages, even non-parallel sentences exhibit similarity scores comparable to those of parallel sentences (see §5.3). This is largely due to the anisotropy problem observed in Transformer models, which can lead to so-called hubness issues, making it difficult to distinguish between similar and dissimilar embeddings (Ethayarajh, 2019), especially in multilingual models (Hämmerl et al., 2023; Rajaee & Pilehvar, 2022). However, a direct comparative analysis of the cosine similarity between parallel and non-parallel sentence pairs across languages can help overcome these issues. Instead of using the absolute cosine similarity value for alignment, we assign binary values (1 or 0) based on whether a criterion for semantic similarity is satisfied. This criterion imposes that (a) parallel sentences should have high cosine similarity, and (b) non-parallel pairs should also have low similarity values, ensuring the similarity is not random or biased. Specifically, if the cosine similarity for a pair of parallel sentences is higher than for any non-parallel sentences, we assign a value of 1 for this pair; otherwise, a value of 0. This approach sidesteps the hubness problem since the absolute cosine similarity values themselves are not directly used.
To compute Mexa, we first apply the cosine similarity function to the pairs of embeddings of parallel sentences in languages and . In Section 3.1, we describe how embeddings can be computed for each layer of the autoregressive language model . We generate a square matrix representing cosine similarities of embeddings at the output of layer for all parallel sentences in languages and . We denote the element in the -th row and -th column of . It represents the cosine similarity between the -th sentence of and the -th sentence of at layer of language model . The diagonal elements of , denoted , represent the cosine similarity between parallel sentence pairs from and . We define the Mexa alignment score () as follows:
@@ 3.1 Sentence Embeddings
Sentence embeddings are a transformation of a sentence into a fixed-size vector that captures its meaning. The process of computing sentence embeddings can vary depending on the model architecture. Typically, sentence embeddings are used in encoder-based models such as BERT, which employ bidirectional attention. In these models, the hidden states of each token are first extracted, then aggregated, commonly by averaging the hidden states from the output layer. Since attention in these models is bidirectional, each token contributes equally to the final embedding. Alternatively one can use the output of the special [CLS] token as per the original BERT work (Devlin et al., 2019).
In this paper, we focus on autoregressive language models that use a decoder-only architecture. In this architecture, attention is not bidirectional; instead, it takes the form of causal attention (left-to-right). In bidirectional attention, each token has access to every other token in the sequence. However, in causal attention, the embedding of a token at position is only influenced by the embedding of preceding tokens at positions . Therefore, simple averaging biases the embeddings towards sentence-initial words. Instead, two alternative methods are considered: using only the last token and weighted averaging. We use and compare both methods to acquire the sentence embeddings needed for Mexa.
A standard way to compute a sentence embedding uses only the last token of that sentence. Jiang et al. (2023b) show that using the last token in the format of a prompt template for a sentence , such as ’This sentence: {s} means in one word:’, can be effective. Inspired by this, Li & Li (2024) used the prompt ’Summarize sentence {s} in one word:’ to obtain the last token embedding as the sentence-level text embedding. However, not all models are instruction-tuned; some earlier works, such as Neelakantan et al. (2022); Wang et al. (2024); Ma et al. (2024), use the last token without any prompt. Since the models studied in this paper are only pre-trained and use multiple languages in the input, we decided to use the last token method without any preceding instruction. An alternative is weighted averaging. It relies on the intuition that using only the last token might not represent the entire sentence, as the influence of earlier tokens may have diminished. This suggests that tokens at the end of the sentence should contribute more to the overall embedding than those at the beginning. Another motivation is that sentence-final tokens are influenced by preceding tokens and contain more context, while the representation of sentence-initial tokens has significantly less contextual representation. To address this, Muennighoff (2022) proposes to assign weights to each token based on its position. Thus, the sentence embedding of layer using position-weighted averaging is:

### v2

@@ Abstract
English-centric large language models (LLMs) often show strong multilingual capabilities. However, their multilingual performance remains unclear and is under-evaluated for many other languages. Most benchmarks for multilinguality focus on classic NLP tasks or cover a minimal number of languages. We introduce MEXA, a method for assessing the multilingual capabilities of pre-trained English-centric LLMs using parallel sentences, which are available for more languages than existing downstream tasks. MEXA leverages that English-centric LLMs use English as a pivot language in their intermediate layers. MEXA computes the alignment between English and non-English languages using parallel sentences to evaluate the transfer of language understanding from English to other languages. This alignment can be used to estimate model performance in different languages. We conduct controlled experiments using various parallel datasets (FLORES-200 and Bible), models (Llama family, Gemma family, Mistral, and OLMo), and established downstream tasks (Belebele, m-MMLU, and m-ARC). We explore different methods to compute embeddings in decoder-only models. Our results show that MEXA, in its default settings, achieves an average Pearson correlation of 0.90 between its predicted scores and actual task performance across languages. This suggests that MEXA is a reliable method for estimating the multilingual capabilities of English-centric LLMs, providing a clearer understanding of their multilingual potential and the inner workings of LLMs.
cis-lmu-mexa.hf.space |
@@
MEXA: Multilingual Evaluation of English-Centric LLMs via Cross-Lingual Alignment
Amir Hossein Kargaran1 Ali Modarressi1 Nafiseh Nikeghbal2 Jana Diesner2 François Yvon3 Hinrich Schütze1 1LMU Munich & Munich Center for Machine Learning 2Technical University of Munich 3Sorbonne Université & CNRS, ISIR {amir, ali}@cis.lmu.de, nafiseh.nikeghbal@tum.de
@@ 1 Introduction
Most state-of-the-art autoregressive large language models (LLMs) are “English-centric”, including closed-source models such as GPT-4 (OpenAI et al., 2023), open-weight models such as Llama 3 (Dubey et al., 2024); and open-source models such as OLMo (Groeneveld et al., 2024). English-centric refers to the majority of the pre-training data for these models being in English (Zhong et al., 2024; Kew et al., 2024). Even models labeled as heavily multilingual, such as BLOOM (BigScience Workshop et al., 2023), have their major pre-training data in English (Laurençon et al., 2022).
@@
Except for open-source models, where pre-training data is available and language distribution is transparent, there is still confusion about the language capabilities and coverage of other LLMs. Primarily, the focus in evaluating LLMs has been on developing benchmarks to assess their performance in English. Most benchmarks in multilingual settings consist of classical monolingual NLP tasks such as sequence labeling (Ahuja et al., 2023; Lai et al., 2023a), automatic translation of popular English benchmarks such as MMLU (Hendrycks et al., 2021) into a limited number of languages (Lai et al., 2023b; OpenAI, 2024), or the creation of language-specific benchmarks (Ghahroodi et al., 2024; Koto et al., 2024; Son et al., 2024; Yüksel et al., 2024; Li et al., 2024c).
@@
Most LLMs are English-centric, either by choice or due to the availability of abundant data sources in English. For these models to be effective in other languages, the other languages must align with the dominant language, i.e., English. Given such alignment, English could act as a “rising tide that raises all ships,” meaning that improvements in English performance could benefit other languages, especially in tasks such as reasoning (Zhu et al., 2024). Contrarily, if a language does not align well with English, an English-centric LLM may not provide meaningful coverage for that language. Indeed, Wendler et al. (2024) have found that for Llama 2 (Touvron et al., 2023b), an English-centric LLM, English could be seen as a kind of “pivot” language, enabling to solve complex semantic tasks in a foreign language through a detour into English. More precisely, they show that Llama 2 was able to decode semantically correct next tokens in the middle layers, assigning higher probabilities to the English tokens than to the foreign version, which is only selected in the upper layers. Zhao et al. (2024) present a hypothesis regarding the middle layers of English-centric LLMs, suggesting that these models use English as a means of reasoning while incorporating multilingual knowledge. Based on their analysis, the number of language-specific neurons in the middle layers decreases within the self-attention mechanism but remains consistent across the layers of the feed-forward structure when processing multilingual queries.
In this paper, we introduce MEXA, a method to estimate the actual multilingual coverage of English-centric LLMs. It builds on the observation that these models semantically use English as a pivot language in their middle layers, by measuring how well embeddings of non-English sentences align with their English counterparts.
@@
We verify the MEXA estimation of language coverage for each LLM, using Pearson correlation between estimated and actual scores for various tasks. We use two parallel datasets: FLORES-200 (NLLB Team et al., 2022) and Bible (Mayer and Cysouw, 2014); nine LLMs: Llama family, Gemma family, Mistral, and OLMo; and three tasks: Belebele (Bandarkar et al., 2024), m-MMLU, and m-ARC (Lai et al., 2023b). Our results show that MEXA achieves a promising average Pearson correlation of 0.90 with established downstream tasks across nine models and two parallel datasets. In our study on the calculation of MEXA scores, we conduct multiple design analyses to examine the impact of token-level pooling for the embeddings (i.e., using the last token versus a weighted average) and layer-level pooling in computing alignment scores. While MEXA demonstrates a high correlation across most setups, we find that a weighted average based on tokens, combined with mean pooling, yields the best results. In summary, MEXA offers a scalable way to estimate the multilingual coverage of English-centric LLMs via alignment with English.
@@ 2 Background and Related Work
We discuss distribution of pre-training data in LLMs, and multilingual evaluation benchmarks in Appendices A.1 and A.2, and focus on cross-lingual alignment here. Research on cross-lingual alignment either aims to uncover the underlying mechanisms of alignment and assesses its impact on models and downstream tasks, or attempts to enhance model performance by enforcing alignment before, during, or after the pre-training phase. Most of these papers have focused on encoder-only models, such as XLM-R (Conneau et al., 2020a) and mBERT (Devlin et al., 2019), among others (Hämmerl et al., 2024). In this work, we focus on decoder-only models.
@@
Understanding Alignment. Ye et al. (2023) show that English-centric models such as Llama 1 (Touvron et al., 2023a) possess multilingual transfer abilities (after fine-tuning on one source language, they can be applied to other languages) and may even surpass the multilingual transfer abilities of multilingual pre-trained models such as BLOOM (BigScience Workshop et al., 2023). Schäfer et al. (2024) find that GPT-2-style decoder-only models show strong cross-lingual generalization when trained on an imbalanced mix of languages. However, when trained on a balanced language set, they do not observe increased performance compared to monolingual settings. Wendler et al. (2024) perform single-token analysis to demonstrate that English-centered LLMs, such as Llama 2, use English semantically as an internal latent language in the middle layers when handling multilingual queries. Zhong et al. (2024) extend this analysis to multiple tokens, also showing that an LLM dominated by both English and Japanese uses both languages as internal latent languages. Zhao et al. (2024) explore how LLMs handle multilingualism. They hypothesize that LLMs initially interpret the query and convert multilingual inputs into English for task-solving. In the middle layers, the models rely on English with self-attention mechanisms for reasoning, while employing multilingual knowledge through feed-forward structures. In the final layers, LLMs generate responses consistent with the original query language. Li et al. (2025) and Li et al. (2024b) are even more closely related to ours. Li et al. (2025) uses absolute cosine similarity values between last token embeddings derived from parallel sentences with English to predict the ranking of language performance across various models. However, as we discuss in Section 3, relying solely on absolute cosine values can be misleading, and as shown in Section 5.3, absolute cosine values are less correlated with downstream tasks than MEXA score. Li et al. (2024b) uses English probing tasks and their automatic translations to construct a multilingual evaluation. While they compare embedding similarity scores between high- and low-resource languages with corresponding evaluation results, similar to Li et al. (2025), they do not assess whether these correlations hold across other downstream tasks. In Section 5, we demonstrate that MEXA scores align closely with a broad range of downstream tasks.
@@
Boosting Alignment. The idea to enforce alignment in encoder-only models using parallel sentences dates back to Conneau and Lample (2019), and has been explored under various guises, e.g., using mixed-language sentences and/or bilingual dictionaries (Huang et al., 2019; Conneau et al., 2020b; Cao et al., 2020; Kulshreshtha et al., 2020; Efimov et al., 2023; Zhang et al., 2023b). Recently, Li et al. (2024d) improved multilingual alignment by initializing the decoder-only models to generate similar representations of aligned words using contrastive learning and preserving this alignment using a code-switching strategy during pre-training. Liu et al. (2024a) propose a data allocation technique to select a core subset of languages for fine-tuning, better aligning the multilingual capabilities of decoder-only LLMs and making them more truthful in their responses. Li et al. (2024a) propose aligning internal sentence representations across different languages using multilingual contrastive learning and aligning outputs by following cross-lingual instructions in the target language for decoder-only models.
@@ 3 MEXA
We now describe the MEXA method for computing the alignment score of language with a pivot language , given the language model . In this paper, we use the term cross-lingual alignment, or simply alignment to refer to the semantic similarity of multilingual embeddings across languages. , for English-centric LLMs and in this paper, is English. To assess alignment, we use parallel sentences in two languages, and . The goal of semantic similarity is to ensure that parallel sentences have sufficiently high similarity, reflecting alignment between the two languages. However, considering only the absolute cosine similarity value as the alignment score does not guarantee proper alignment. For some languages, even non-parallel sentences exhibit similarity scores comparable to those of parallel sentences (see §5.3). This is largely due to the anisotropy problem observed in transformer models, which can lead to so-called hubness issues, making it difficult to distinguish between similar and dissimilar embeddings (Ethayarajh, 2019), especially in multilingual models (Hämmerl et al., 2023; Rajaee and Pilehvar, 2022). However, a direct comparative analysis of the cosine similarity between parallel and non-parallel sentence pairs across languages can help overcome these issues. Instead of using the absolute cosine similarity value for alignment, we assign binary values (1 or 0) based on whether a criterion for semantic similarity is satisfied. Our criterion imposes that (a) parallel sentences should have high cosine similarity, and (b) non-parallel pairs should also have low similarity values, ensuring the similarity is not random or biased. Specifically, if the cosine similarity for a pair of parallel sentences is higher than for any non-parallel sentences, we assign a value of 1 for this pair; otherwise, a value of 0. This approach sidesteps the hubness problem since the absolute cosine similarity values themselves are not directly used.
To compute MEXA, we first apply the cosine similarity function to the pairs of embeddings of parallel sentences in languages and . In Section 3.1, we describe how embeddings can be computed for each layer of the autoregressive language model . We generate a square matrix representing cosine similarities of embeddings at the output of layer for all parallel sentences in languages and . We denote the element in the -th row and -th column of . It represents the cosine similarity between the -th sentence of and the -th sentence of at layer of language model . The diagonal elements of , denoted , represent the cosine similarity between parallel sentence pairs from and . We define the MEXA alignment score as follows:
@@ 3.1 Sentence Embeddings
We focus on autoregressive language models that use a decoder-only architecture. In this architecture, attention is not bidirectional; instead, it takes the form of causal attention (left-to-right). In bidirectional attention, each token has access to every other token in the sequence. However, in causal attention, the embedding of a token at position is only influenced by the embedding of preceding tokens at positions . Therefore, simple averaging values biases the embeddings towards sentence-initial words. Instead, we consider alternative methods: using only the last token and weighted averaging. We use and compare both methods to acquire the sentence embeddings needed for MEXA.
A standard way to compute a sentence embedding uses only the last token of that sentence. Jiang et al. (2023b) show that using the last token in the format of a prompt template for a sentence , such as ’This sentence: {s} means in one word:’, can be effective. Inspired by this, Li and Li (2024) used the prompt ’Summarize sentence {s} in one word:’ to obtain the last token embedding as the sentence-level text embedding. However, not all models are instruction-tuned; some earlier works, such as Neelakantan et al. (2022); Wang et al. (2024); Ma et al. (2024), use the last token without any prompt. Since the models studied in this paper are only pre-trained and use multiple languages in the input, we decided to use the last token method without any preceding instruction. An alternative is weighted averaging, which relies on the intuition that using only the last token might not represent the entire sentence, as the influence of earlier tokens may have diminished. This implies that the tokens at the end of the sentence should contribute more to the overall embedding than those at the beginning. Another motivation for using weighted averaging is that sentence-final tokens are influenced by preceding tokens and contain more context, while the representation of sentence-initial tokens has significantly less contextual representation. To address this, Muennighoff (2022) proposes to assign weights to each token based on its position. Thus, the sentence embedding of layer using position-weighted averaging is:
@@
3.2 Robustness of MEXA
We show that the MEXA alignment score () is very robust, and the odds of this score randomly achieving a high value are very slim. Recall that measures the fraction of diagonal elements in matrix that have the maximum value in their respective rows and columns. If this condition is met times out of diagonal elements, then is . In an random matrix, the probability of a diagonal element being the maximum in its row and column (a total of elements) is . The probability that at least out of independent variables are satisfied, given that the diagonal element is the maximum in its row and column, can be computed using the binomial distribution: